# ClusterRCA: Network Failure Diagnosis in HPC Systems Using Multimodal Data

### v1

@@ III-C1 State classifier
Classifier-based diagnosis methods distinguish root causes from other states by using well-trained classifiers, which is also useful in our approach. We previously extracted node metrics and logs as features, but the combinations were too complex for effective classification. Therefore, we first apply anomaly detection for feature compression. As observed, nodes with similar traffic patterns tend to have similar characteristics, but anomalies alter a few. We compare features to normal nodes, capture anomalies, and use anomaly vectors to train the classifier. For feature vectors to classify, we evaluate the similarity between the vector and normal samples using Formula 1, where $V_{a}$ is node $a$’s feature vector and $M$ is the dimension.
$similarity\left(V_{a},V_{Normal}\right)=\frac{\sum_{M}^{i=1}\mathbb{I}\left(V_{a}\left[i\right]=V_{Normal}\left[i\right]\right)}{M}$ (1)
Since most features of the most similar normal sample match the sample, they can be considered in a very similar state. Features that differ between the sample and normal sample are labeled anomalies (1). Matching features are normal (0). We transform the discrete feature vector into a 0-1 vector, simplifying the multi-dimensional feature space into a compressed representation of the anomalies that differentiate the sample from normal.
Figure 6: The training and predicting of state classifier.
As Fig. 6 shows, we train a status classifier on historical NIC pair feature vectors and operator-provided labels (nodes with various types of labels in Figure  reffig:classifier are all NIC pairs). Operators label only the culprit node and failure type for each historical failure. Other nodes are automatically labeled as victims. Each labeling provides one culprit node and multiple victim nodes. The well-trained classifier predicts a state probability matrix for a realtime feature vector, specifying the probabilities of nodes being the culprit of various failure types, victim, or normal. The multi-classifier used in this paper is the random forest [33], which has been proven to outperform other classifiers in failure diagnosis for HPC system [18]. However, the classifier’s accuracy still needs to be improved to determine the precise failure type conclusively or identify the culprit node uniquely, as there are typically multiple nodes exhibiting high probabilities of different failure types.
@@ III-C2 Customized Random Walk
  • Culprit state: Node $N_{i}$ is in the culprit state with probability $PF_{Ni}=P_{Ni}[F_{1}]+P_{Ni}[F_{2}]+...+P_{Ni}[F_{k}]$. If the walker is at the culprit node, it should remain as this is the root cause.
@@
  • Normal state: Nodes in normal states show little anomalies, but the network may not be healthy. We want the walker to quickly leave these nodes, randomly selecting neighbors. The probability is $P_{Ni}[Normal]/|B_{Ni}|$, where $|B_{Ni}|$ signifies the number of $N_{i}$ neighbors.
Based on the above three states, we define the transition probability matrix $Q$. Specifically, the probability of the current node $N_{i}$ transitioning to adjacent node $N_{j}$ can be calculated as follows:
$Q_{ij}=\left\{\begin{aligned} PF_{i},\qquad\qquad\qquad\qquad&if\ j=i\\ \left(\begin{aligned} P_{Ni}[Victim]\times\frac{PF_{Nj}}{{\textstyle\sum_{k\in B_{Ni}}PF_{Nk}}}\\ +\frac{P_{Ni}[Normal]}{|B_{Ni}|}\qquad\quad\end{aligned}\right),&if\ j\neq i\ and\ N_{j}\in B_{Ni}\\ \end{aligned}\right.$ (2)
We then normalize each row of $Q_{ij}$ to obtain the final transition probability matrix. Each row of the matrix represents the probability of the walker transitioning from one node to other nodes in the graph, with the sum of these probabilities equaling 1. We can perform the rapid random walk with the transition probability matrix, as shown in Algorithm 1.
@@
input : number of results: $M$, number of walking steps: $STEPS$, list of states probability matrix: $S$, number of nodes: $N$
output : A ordered list of root cause: $result$
@@
3Initialize $result\leftarrow[]$, start node $V\leftarrow$ RandomInt($N$);
@@ 5 Initialize transition probability matrix $Q\leftarrow$ GetTransitionProbabilityMatrix($S$);
6
7for $i\leftarrow 1$ to $M$ do
8       for $j\leftarrow 1$ to $STEPS$ do
9             $V\leftarrow$ ProbabilityBasedRandomChoose($N$, $Q[V]$);
10             COUNTS $[V]$leftarrow COUNTS $[V]$ + $1$;
11
@@ 13      $nodeId\leftarrow$ index of Max(COUNTS);
[…]
15       add $(nodeId,faultId)$ to $result$;
@@ 20       COUNTS $\leftarrow[0]*N$;
The algorithm initiates a random walk from an arbitrarily chosen initial node, traversing between nodes according to the transition probability matrix. It then determines the culprit node by checking the number of visits to each node. When the number of transitions exceeds the specified count $STEP$, the most visited node is the diagnostic culprit node. Its failure type with the highest state probability is the diagnostic failure type. The diagnostic result, which identifies the culprit node and failure type, is added to $result$. The algorithm converts the probability of the corresponding failure type for the last diagnostic culprit node to its victim state probability to obtain more combinations of culprit nodes and failure types. It then updates the transition probability matrix for the next iteration. After $M$ iterations, the algorithm returns an ordered list of $M$ root causes, each specifying the culprit node and associated failure type. The approach uses a random walk algorithm to find the most likely culprit node when classification alone cannot provide a definitive conclusion due to limited accuracy.
@@ IV-A1 Dataset
To evaluate the performance of ClusterRCA, we establish a small-scale HPC computing cluster environment in a top-tier global HPC device vendor ($\boldsymbol{A}$ company)’s laboratory. The environment comprises four TaiShan server computing nodes and one network switch node. Each computing node connects to the switch node through a high-speed 100Gbps network. We deploy multiple HPC applications within this experimental setup and simulate network failures using script injection. Throughout the execution of the HPC applications, we continuously collect monitoring data, including metrics, logs, and topology information from the computing and switch nodes.
@@
Based on these HPC applications, we collect six datasets, including 739 network failure samples and 117 normal samples. We collect 176 network interface metrics and 38 computing resource metrics for each computing node and connected switch node port for each sample. The total number of log entries is 105,095. Each dataset contains samples of the failure types mentioned in Section II, namely F1: wrong packet, F2: PFC storm, F3: switch port performance restricted, F4: Tx timeout, F5: PFC parameter unaligned, F6: switch PFC disabled and F7: computing node PFC disabled. The samples for various failure types in the datasets are listed in Table II. It is worth noting that due to the unique traffic characteristics of LAMMPS, it is almost unaffected by certain flow control error-related failures (no network congestion occurred). Consequently, we cannot collect valid samples for such failures in this application scenario. Furthermore, since the training process of ClusterRCA requires normal NIC-pair samples, we further divided the samples from the WRF[26] application into WRF_A and WRF_B. WRF_A contains all normal NIC-pair samples and approximately half of the failure NIC-pair samples used for training ClusterRCA, while WRF_B is used to evaluate the failure diagnosis performance of ClusterRCA.
@@
Dataset NIC failures Link bottlenecks Flow control failures Normal Total
@@ IV-A2 Baseline methods
Due to the different formats of the analysis results of these methods, we only discuss whether their results are accurate during the experiment. Specifically, for AutoMAP [19] and Cloud19 [20], we only access the accuracy of their culprit node localization. Conversely, the evaluation of LogCluster [22] and CloudRCA [3] concentrates on determining the accuracy of failure type identification. For ClusterRCA and OnlineDiagnosis[18], our evaluation encompasses accurately assessing identified culprit nodes and failure types.
@@ IV-A3 Evaluation metrics
${RC_{i}}_{a}$ is the actual root cause of test case $a$ (culprit node or failure type or both), and ${RC_{s}}_{a}k$ is the failure diagnosis result list of the diagnosis method.
@@ IV-B Overall Performance (RQ1)
We first use WRF_A as the training set and WRF_B as the testing set to evaluate the overall performance of ClusterRCA and baselines. The two sets were collected when the same HPC application (WRF[26]) was running. The accuracy of ClusterRCA and the baseline methods are listed in Table III. Notably, the failure diagnosis capability of ClusterRCA significantly surpasses other methods, achieving a top-1 accuracy rate of 0.98. Cloud19[20], limited by its search space, can only enumerate four root causes, resulting in a considerably lower $AC@1$ compared to ClusterRCA. LogCluster[22] solely considers log data, its $AC@1$ is limited to 0.20. In contrast, CloudRCA[3] integrates information from logs and metrics data and achieves higher accuracy. OnlineDiagnosis[18] achieves considerable effectiveness ($Avg@5$ $>$ 0.95) because the statistical features extracted by OnlineDiagnosis accurately distinguish nodes’ various network flow sizes. However, its accuracy will significantly decrease when the HPC application scenarios are inconsistent in the training and test sets.
@@ IV-C Robustness Evaluation (RQ2)
In practice, the applications running on HPC systems are diverse and constantly evolving, and the failure pattern caused by network failures also changes accordingly. For system administrators, running HPC applications is an uncontrollable user behavior. Therefore, we will analyze how ClusterRCA maintains performance across different HPC applications to determine whether it can adapt to the real network environment. Specifically, we use the WRF_A to train ClusterRCA and baseline methods. Then, we use test case sets of other applications except WRF[26] to test these methods’ failure diagnosis accuracy under different HPC applications. Cloud19[20], offering only four root cause candidates, is excluded from this evaluation due to the unavailability of $Avg@5$ calculation. The test case sets used for testing include Grapes[31], QE[27], GROMACS[29], LAMMPS[28], and OpenFOAM[30], and the results are shown in Fig 7.
The first five subgraphs show that ClusterRCA can achieve $AC@1$ above 0.9 in five HPC applications, which is extremely higher than the baseline methods. The subgraph of $Avg@5$ shows that ClusterRCA has a significant failure diagnosis accuracy as well. Although OnlineDiagnosis[18] achieves an $Avg@5$ of 0.95 in the overall performance experiment, due to the fact that the data feature extraction effect of OnlineDiagnosis[18] is more linked to the application data, under different HPC applications, its $Avg@5$ is only at most 0.73. Also, it can be observed that CloudRCA[3] can reach the level of ClusterRCA in terms of Top5 accuracy because of more data features considered, but the results of the top root cause ranking are very bad. Both AutoMAP[19] and LogCluster[22], on the other hand, are limited in their accuracy due to the fact that they only consider one data modality (metric or log data).
@@ IV-D Ablation Study (RQ3)
In the ablation experiments, this work systematically removed some components of ClusterRCA and evaluated the performance of the remaining parts. By comparing the performance of the complete method and the ablation schemes, we can better understand the contribution of each component to ClusterRCA. First, we individually eliminated the features and data in the feature engineering part. Among them, we extracted two types of features for the metrics data: pattern features and level features. We deleted one of them respectively and consequently obtained two variants of ClusterRCA. C1: Remove metrics pattern features; C2: Remove metrics mean level features. Then, we eliminated the input data. C3: Remove metric data input; C4: Remove log data input. Finally, we consider the random walk components. C5: Remove the random walk component. Deleting the entire state classifier component would cause the system to be unable to yield the failure diagnosis result list, which has not yet been tested.
@@
Fig 8 shows the experimental results of ClusterRCA and its variants, using the samples of WRF_A for training and all our HPC applications for testing. ClusterRCA outperforms all variants in these tests. From subgraph $Avg@5$ (WRF_B), it is apparent that the lack of input data (C3 & C4) significantly reduces the learning ability of the system, resulting in $Avg@5$ during training and testing on the same application (WRF[26]) is much lower than that of ClusterRCA and other variants. The subgraph of $Avg@5$ (Others) shows the distribution of $Avg@5$ of ClusterRCA and variants across non-training application sample sets and marks the mean of $Avg@5$. Among them, the removal of some metrics feature types (C1 & C2) has a lower $Avg@5$ (0.90 & 0.96) in other HPC applications than ClusterRCA (0.98), but not as obvious as directly removing the metrics data (C3, 0.66). It indicates that sufficient data types are more important for ClusterRCA than abundant feature types. Comparing the average $Avg@5$ of ClusterRCA (0.98) and C5 (0.90), it is evident that combining the random walk can achieve better diagnostic effectiveness than using the classifier alone.
@@
Overall, each component contributes positively to the final diagnostic performance, and together they enable ClusterRCA to achieve the best possible results. Input data (metric and log data) are the foundation of the model’s ability to learn system behavioral patterns; removing either data causes the model to lose the ability to sense important information, especially metrics data, which is a direct reflection of the system’s performance and health. Feature engineering, helps the model capture failure-related patterns more effectively by distilling key information from the raw data; removing some feature types will affect the integrity of the information. The random walk component captures dynamic correlations or shifting patterns between different data sources thus improving diagnostic accuracy.

### v2

@@ III-C1 State classifier
Classifier-based diagnosis methods distinguish root causes from other states using well-trained classifiers, which is also helpful in our approach. We previously extracted node metrics and logs as features, but the combinations were too complex for effective classification. Therefore, we first apply anomaly detection for feature compression.
Our observation that nodes with similar traffic patterns exhibit similar characteristics, with anomalies altering a few, is rooted in the inherent design of HPC systems. Identically configured nodes executing similar workloads naturally produce highly consistent operational behaviors, reflected in their performance metrics and log events. Anomalies, by definition, represent deviations from this expected consistency, thereby altering a subset of these features.
To capture features that are shifted due to anomalies, we compare features to normal nodes. We maintain a feature vector sample library that does not contain abnormal jobs and calculate the similarity between the real-time feature vectors and each normal sample in the library using Formula 1, where $V_{r}$ is the real-time feature vector, $M$ is the dimension and $V_{Normal}$ is the feature vectors in the normal sample library.
$similarity\left(V_{r},V_{Normal}\right)=\frac{\sum_{M}^{i=1}\mathbb{I}\left(V_{r}\left[i\right]=V_{Normal}\left[i\right]\right)}{M}$ (1)
Since most features of the most similar normal sample match those of the sample, they can be considered to be in a very similar state. We compare the current sample with the normal sample position by position. Positions where there are differences between the sample and the normal sample are marked as abnormal values “1”, and positions where there are the same are marked as normal values “0”. In this way, the feature vector is converted into a 0-1 vector of the same dimension, simplifying the multi-dimensional feature space into a compressed representation of the anomalies that differentiate the sample from normal. In addition, if the current work is extended to another HPC network, it is only necessary to add feature vectors that represent typical data patterns to the normal sample library, without the need for retraining. Subsequently, this 0-1 vector is input into the state classifier.
Figure 6: The training and predicting of the state classifier.
When performing fault injection, operators simulate a network failure on a specific node within a job, which is defined as the “culprit” node. Other nodes under this job are labeled as “victim” nodes, while the remaining nodes are classified as “normal” nodes.
As Fig. 6 shows, we train a status classifier on historical NIC pair feature vectors and operator-provided labels (nodes with various types of labels in Fig. 6 are all NIC pairs). The well-trained classifier predicts a state probability matrix for a real-time feature vector, specifying the probabilities of nodes being the culprit of various failure types, a victim, or normal.
The multi-classifier used in this paper is the random forest [32], which has been proven to outperform other classifiers in failure diagnosis for HPC systems [6]. However, the results based on the classifier still need further analysis to determine the precise failure type conclusively or identify the culprit node uniquely, as there are typically multiple nodes exhibiting high probabilities of different failure types. Therefore, based on the preliminary results of the state classifier, we still need to use graph-based methods to determine the location of the root cause node.
@@ III-C2 Customized Random Walk
  • Culprit state: Node $N_{i}$ is in the culprit state with probability $PF_{Ni}=P_{Ni}[F_{1}]+P_{Ni}[F_{2}]+...+P_{Ni}[F_{k}]$. If the walker is at the culprit node, it should remain, as this is the root cause.
@@
  • Normal state: Nodes in normal states show few anomalies, but the network may not be healthy. We want the walker to leave these nodes, randomly selecting neighbors quickly. The probability is $P_{Ni}[Normal]/|B_{Ni}|$, where $|B_{Ni}|$ signifies the number of $N_{i}$ neighbors.
Based on the above three states, we define the transition probability matrix $Q$. Specifically, the probability of the current node $N_{i}$ transitioning to the adjacent node $N_{j}$ can be calculated as follows:
[…]
We then normalize each row of $Q_{ij}$ to obtain the final transition probability matrix. Each row of the matrix represents the probability of the walker transitioning from one node to another node in the graph, with the sum of these probabilities equaling 1. We can perform the rapid random walk with the transition probability matrix, as shown in Algorithm 1.
@@
input : number of results: $M$, number of walking steps: STEPSSTEPS, list of states probability matrix: $S$, number of nodes: NN
output : A ordered list of root cause: resultresult
@@
31ex
4Initialize $result\leftarrow[]$, start node $V\leftarrow$ RandomInt(NN);
@@ 6 Initialize transition probability matrix $Q\leftarrow$ GetTransitionProbabilityMatrix($S$);
7
8for $i\leftarrow 1$ to $M$ do
9  for $j\leftarrow 1$ to STEPSSTEPS do
10     $V\leftarrow$ ProbabilityBasedRandomChoose(NN, $Q[V]$);
11     COUNTS $[V]$ $\leftarrow$ COUNTS $[V]$ + 11;
12
@@ 14 $nodeId\leftarrow$ index of Max(COUNTS);
15  $faultId\leftarrow$ index of Max($S[nodeId][F_{0}:F_{k}]$);
16  add $(nodeId,faultId)$ to resultresult;
@@ 21  COUNTS $\leftarrow[0]*N$;
The algorithm initiates a random walk from an arbitrarily chosen initial node, traversing between nodes according to the transition probability matrix. It then determines the culprit node by checking the number of visits to each node. When the number of transitions exceeds the specified count $STEP$, the most visited node is the diagnostic culprit node. Its failure type with the highest state probability is the diagnostic failure type. The diagnostic result, which identifies the culprit node and failure type, is added to resultresult.
Since the granularity of random walks is node-level, the final results must be refined to a more specific fault type level. If the root cause list is output based solely on the results of a single walk, it may overlook situations where multiple faults occur on a single node. Therefore, the algorithm converts the probability of the corresponding failure type for the last diagnostic culprit node to its victim state probability to obtain more combinations of culprit nodes and failure types. It then updates the transition probability matrix for the next iteration. After $M$ iterations, the algorithm returns an ordered list of $M$ root causes, each specifying the culprit node and associated failure type.
@@ IV-A1 Dataset
To evaluate the performance of ClusterRCA, we establish a small-scale HPC computing cluster environment in a top-tier global HPC device vendor ($\boldsymbol{A}$ company)’s laboratory. The environment comprises four TaiShan server computing nodes and one network switch node. Each computing node connects to the switch node through a high-speed 100Gbps network. We deploy multiple HPC applications within this experimental setup and simulate a variety of network failures.
@@
Based on these HPC applications, we collect six datasets, including 739 network failure samples and 117 normal samples. We collect 176 network interface metrics and 38 computing resource metrics for each computing node and connected switch node port for each sample. The total number of log entries is 105,095. Each dataset contains samples of the failure types mentioned in Section II, namely F1: wrong packet, F2: PFC storm, F3: switch port performance restricted, F4: Tx timeout, F5: PFC parameter unaligned, F6: switch PFC disabled, and F7: computing node PFC disabled.
These network failures are simulated primarily through the injection of dedicated scripts. This approach allowed us to alter system-level parameters dynamically, reconfigure network device settings, and leverage specialized firmware utilities (e.g., Mellanox tools) to induce various fault mechanisms, such as error packets or periodic packet loss. Furthermore, we utilized packet generation tools and direct system calls to trigger specific queue behaviors, including Tx Timeout by stopping sub-queues. Throughout the execution of the HPC applications, we continuously collect monitoring data, including metrics, logs, and topology information from the computing and switch nodes.
The samples for various failure types in the datasets are listed in Table II. It is worth noting that due to the unique traffic characteristics of LAMMPS, it is almost unaffected by certain flow control error-related failures (no network congestion occurred). Consequently, we cannot collect valid samples for such failures in this application scenario. Furthermore, since the training process of ClusterRCA requires normal NIC-pair samples, we further divided the samples from the WRF[37] application into WRF_A and WRF_B. WRF_A contains all normal NIC-pair samples and approximately half of the failure NIC-pair samples used for training ClusterRCA, while WRF_B is used to evaluate the failure diagnosis performance of ClusterRCA.
@@
Dataset
NIC
failures
Link
bottlenecks
Flow control
failures
Normal Total
@@ IV-A2 Baseline methods
Due to the different formats of these methods’ analysis results, we only discuss the accuracy of the results they can output in the experiment. Specifically, we only access the accuracy of their culprit node localization for AutoMAP [12] and Cloud19 [7]. Conversely, the evaluation of LogCluster [9] and CloudRCA [14] concentrates on determining the accuracy of failure type identification. For ClusterRCA and OnlineDiagnosis[6], our evaluation encompasses accurately assessing identified culprit nodes and failure types.
@@ IV-A3 Evaluation metrics
${RC_{i}}_{a}$ is the actual root cause of test case $a$ (culprit node or failure type or both), and ${RC_{s}}_{a}k$ is the first $k$ results in the failure diagnosis result list of the diagnosis method.
@@ IV-B Overall Performance (RQ1)
We first use WRF_A as the training set and WRF_B as the testing set to evaluate the overall performance of ClusterRCA and baselines. The two sets were collected when the same HPC application (WRF[37]) was running. The accuracy of ClusterRCA and the baseline methods are listed in Table III. Notably, the failure diagnosis capability of ClusterRCA significantly surpasses other methods, achieving a top-1 accuracy rate of 0.98. Cloud19[7], limited by its search space, can only enumerate four root causes, resulting in a considerably lower $AC@1$ compared to ClusterRCA. LogCluster[9] solely considers log data; its $AC@1$ is limited to 0.20. In contrast, CloudRCA[14] integrates information from logs and metrics data and achieves higher accuracy. OnlineDiagnosis[6] achieves considerable effectiveness ($Avg@5$ >> 0.95) because the statistical features extracted by OnlineDiagnosis accurately distinguish nodes’ various network flow sizes. However, its accuracy will significantly decrease when the HPC application scenarios are inconsistent in the training and test sets.
@@ IV-C Robustness Evaluation (RQ2)
In practice, the applications running on HPC systems are diverse and constantly evolving, and the failure pattern caused by network failures also changes accordingly. For system administrators, running HPC applications is an uncontrollable user behavior. Therefore, we will analyze how ClusterRCA maintains performance across different HPC applications to determine whether it can adapt to the real network environment. Specifically, we use the WRF_A to train ClusterRCA and baseline methods. Then, we use test case sets of other applications except WRF[37] to test these methods’ failure diagnosis accuracy under different HPC applications. Cloud19[7], offering only four root cause candidates, is excluded from this evaluation due to the unavailability of the $Avg@5$ calculation. The test case sets used for testing include Grapes[38], QE[39], GROMACS[40], LAMMPS[41], and OpenFOAM[42], and the results are shown in Fig. 7.
The first five subgraphs show that ClusterRCA can achieve $AC@1$ above 0.9 in five HPC applications, significantly higher than the baseline methods. The subgraph of $Avg@5$ shows that ClusterRCA has a significant failure diagnosis accuracy. Although OnlineDiagnosis[6] achieves an $Avg@5$ of 0.95 in the overall performance experiment, because the data feature extraction effect of OnlineDiagnosis[6] is more linked to the application data, under different HPC applications, its $Avg@5$ is only at most 0.73. Also, it can be observed that CloudRCA[14] can reach the level of ClusterRCA in terms of Top5 accuracy because of more data features considered, but the results of the top root cause ranking are terrible. Both AutoMAP[12] and LogCluster[9], on the other hand, are limited in their accuracy because that they only consider one data modality (metric or log data).
@@ IV-D Ablation Study (RQ3)
In the ablation experiments, this work systematically removed some components of ClusterRCA and evaluated the performance of the remaining parts. By comparing the performance of the complete method and the ablation schemes, we can better understand the contribution of each component to ClusterRCA. First, we individually eliminated the features and data in the feature engineering part. Among them, we extracted two types of features for the metrics data: pattern features and level features. We deleted one of them respectively and consequently obtained two variants of ClusterRCA. Specifically, C1: Remove metrics pattern features; C2: Remove metrics mean level features. Then, we eliminated the input data. C3: Remove metric data input; C4: Remove log data input. Finally, we consider the random walk components. C5: Remove the random walk component. Deleting the entire state classifier component would cause the system to be unable to yield the failure diagnosis result list, which has not yet been tested.
@@
Fig. 8 shows the experimental results of ClusterRCA and its variants, using the samples of WRF_A for training and all our HPC applications for testing. ClusterRCA outperforms all variants in these tests. From subgraph $Avg@5$ (WRF_B), it is apparent that the lack of input data (C3 & C4) significantly reduces the learning ability of the system, resulting in $Avg@5$ during training and testing on the same application (WRF[37]) is much lower than that of ClusterRCA and other variants. Also, since different HPC applications have different traffic patterns and communication behaviors, the importance of metrics and logs varies. Therefore, in some applications, metrics alone perform better than logs alone, while in others, it is the reverse. The subgraph of $Avg@5$ (Others) shows the distribution of $Avg@5$ of ClusterRCA and variants across non-training application sample sets and marks the mean of $Avg@5$. Among them, the removal of some metrics feature types (C1 & C2) has a lower $Avg@5$ (0.90 & 0.96) in other HPC applications than ClusterRCA (0.98), but not as apparent as directly removing the metrics data (C3, 0.66). It indicates that sufficient data types are more critical for ClusterRCA than abundant feature types. Comparing the average $Avg@5$ of ClusterRCA (0.98) and C5 (0.90), it is evident that combining the random walk can achieve better diagnostic effectiveness than using the classifier alone.
@@
Overall, each component contributes positively to the final diagnostic performance, and together they enable ClusterRCA to achieve the best possible results. Metric and log data are the foundation of the model’s ability to learn system behavioral patterns; removing either data causes the model to lose the ability to sense important information, especially metric data, which directly reflects the system’s performance and health. Feature engineering helps the model capture failure-related patterns more effectively by distilling key information from the raw data; removing some feature types will affect the integrity of the information. The random walk component captures dynamic correlations or shifting patterns between different data sources, thus improving diagnostic accuracy.